# Group Distributionally Robust Dataset Distillation with Risk Minimization

### v2

@@ Abstract
Dataset distillation (DD) has emerged as a widely adopted technique for crafting a synthetic dataset that captures the essential information of a training dataset, facilitating the training of accurate neural models. Its applications span various domains, including transfer learning, federated learning, and neural architecture search. The most popular methods for constructing the synthetic data rely on matching the convergence properties of training the model with the synthetic dataset and the training dataset. However, targeting the training dataset must be thought of as auxiliary in the same sense that the training set is an approximate substitute for the population distribution, and the latter is the data of interest. Yet despite its popularity, an aspect that remains unexplored is the relationship of DD to its generalization, particularly across uncommon subgroups. That is, how can we ensure that a model trained on the synthetic dataset performs well when faced with samples from regions with low population density? Here, the representativeness and coverage of the dataset become salient over the guaranteed training error at inference. Drawing inspiration from distributionally robust optimization, we introduce an algorithm that combines clustering with the minimization of a risk measure on the loss to conduct DD. We provide a theoretical rationale for our approach and demonstrate its effective generalization and robustness across subgroups through numerical experiments. Source Code.
Machine Learning, ICML
@@ 1 Introduction
Dataset distillation (DD) is a burgeoning area of interest, involving the creation of a synthetic dataset significantly smaller than a training set yet demonstrating comparable performance on a model. This practice has gained prominence in various computation-sensitive applications, providing a valuable means to efficiently construct accurate models. The standard optimization objectives that are used to steer the construction of the synthetic data typically aim to foster either distributional similarity to the training set or exhibit similar stochastic gradient descent (SGD) training dynamics as the original dataset. Notably, recent literature suggests that the latter category has proven more successful. Intuitively, this success can be attributed to the rationale that, with considerably fewer samples, prioritizing the most relevant information for training and model building becomes more judicious.
However, in practical terms, executing the same training task with a reduced computational load is not universally applicable, nor is it the standard application of DD. Instead, the usual scenario involves re-training and inference on a synthetic dataset under circumstances that are similar but perturbed to the original task of interest. To this end, we aim to execute DD in a manner that specifically addresses the challenges posed by distributional shifts. Additionally, in the context of online inference, it is common to observe alterations in the overall prior distribution of samples, particularly in terms of subgroups. In practice, within a stream of samples, there are often clumps belonging to specific subgroups, collectively representing a small portion of the sample space. The ability of DD to facilitate the construction of a model that performs effectively across time varying distributional shifts and subgroup appearance is a noteworthy practical concern. This aspect has yet to be prioritized in DD literature, as per our current understanding.
@@
When evaluating test accuracy, a commonly utilized algorithmic framework is Distributionally Robust Optimization (DRO). Within DRO, the approach involves solving a bilevel optimization problem where the objective is to minimize the loss over a data distribution, subject to the constraint that this distribution is a specific probabilistic distance away from the population distribution:
$\min\limits_{\theta}\max\limits_{\mathcal{Q}\in\mathbb{Q}},\,F(\theta,\mathcal{Q}),\,\mathbb{Q}=\{\mathcal{Q}:\mathcal{D}(\mathcal{Q}||\mathcal{P})\leq\Delta\}$ (1)
@@
where $F$ is the loss function, $\theta$ are the parameters, $\mathbb{Q}$ is the distributional uncertainty set, $\mathcal{D}$ is an f-divergence and $\mathcal{P}$ is the population distiribution.
Intuitively, there are two considerations through which DRO can be understood to be the appropriate framework for addressing the generalization concerns discussed above. First, one can consider that a test and training distribution are both subsamples from the same population distribution, and so there must be a significant resemblance between the two. Indeed, the work (Xu & Mannor, 2012) presented arguments indicating theoretical equivalence of testing to training error as robustness to perturbations of the data distribution, and the generalization accuracy.
@@
Second, a conventional least likelihood classifier inherently assigns greater weight to samples with larger prior distributions compared to those with lower overall probability density. This inherent bias poses a fundamental risk of underperformance in small subgroups within the population. DRO has been conjectured and observed to mitigate this, both generically (Duchi & Namkoong, 2021) and with explicit quantification of subgroups in the distributional uncertainty set (Oren et al., 2019).
Nonetheless, the conventional formulation of DRO, as depicted in (1), has been revealed to be excessively conservative, effectively optimizing towards the exact training distribution and hindering practical generalization (Hu et al., 2018). Research indicates that a more precisely defined distributional uncertainty set, focusing on specific collections of probability distributions from subsamples of the population rather than encompassing all possibilities in a ball, yields improved generalization performance. An illustrative example of this refined approach to DRO can be found in (Oren et al., 2019), where subgroups of topics are considered for training language models.
@@
Figure 1: A visual representation of the robust inference task involves the partial partitioning of the population distribution across subgroups. A classifier is considered robust when it demonstrates high performance regardless of the specific group’s data it is conditioned on. For instance, at a particular time, a steady stream of samples from $\mathcal{X}|c_{3}$ may appear to the classifier, while it may be that the region of sample space defined by this subgroup is both geometrically small (as in the Figure) as well as having low overall prior density. If this subgroup’s behavior is particularly anomalous, a model trained only on minimizing empirical risk may perform poorly on this subgroup, and a synthetic dataset aiming to ensure similar training will poorly represent $\mathcal{X}|c_{3}$.
Inspired by these concepts, this paper introduces a DD procedure designed specifically for this objective. We synchronize the dataset construction with the loss measured by robust criteria. To circumvent the challenges posed by the intractable bilevel DRO, we employ risk measures applied to the data loss as the criteria during training. At each iteration, a zero-order gradient estimate of the synthetic dataset, considering the sampled data’s performance on our robust criteria, is calculated. The bilevel DD problem is then minimized by a sequence of these iterations.
This presents an important shift in perspective with respect to dataset distillation. Standard consideration of DD broadly considers each data point independently, rather than their interrelationships. This is unfortunate, as the small size of the synthetic dataset suggests that there is both tractability in computing collective valuations of the data, and necessity in ensuring parsimony with minimal redundancy in the information in the dataset. In the context of deploying memory-sensitive AI pipelines for DD, one can anticipate the implementation of time-ordered stratified sampling across subgroups and adaptations to address general distributional shifts. Giving thoughtful attention to the dataset’s coverage across the actual support can markedly enhance the practical utility of DD.
@@
Our paper continues below as follows. In Section 2 we review related works. Next, we present our Algorithm, we analyze the optimization problem defining the notion of generalization of interest and argue how our method effectively targets this criterion in Section 3. In Section 4 we present Numerical Results validating the superior overall and subgroup generalization performance on standard machine learning benchmarks. Finally, we conclude this work in Section 5.
2 Related Work
2.1 Dataset Distillation
Dataset distillation (DD) seeks to distill the richness of extensive datasets into compact sets of synthetic images that closely mimic training performance. These condensed images prove invaluable for various tasks, including continual learning (Gu et al., 2023b), federated learning (Liu et al., 2022; Jia et al., 2023), neural architecture search (Such et al., 2020), and semi-supervised learning (Vahidian et al., 2020; Joneidi et al., 2020). Existing DD methodologies can be broadly categorized into bi-level optimization and training metric matching approaches. Bi-level optimization integrates meta-learning into the surrogate image update process (Zhou et al., 2022; Loo et al., 2023). Conversely, metric matching techniques refine synthetic images by aligning with training gradients (Kim et al., 2022; Liu et al., 2023), feature distribution (Sajedi et al., 2023; Zhao et al., 2023), or training trajectories (Wu et al., 2023; Du et al., 2023) compared to the original images. There are also some methods involving generative prior for DD (Cazenavette et al., 2022; Wang et al., 2023a; Gu et al., 2023a).
2.2 Robustness in Dataset Distillation
To the best of our knowledge, this is the first work on subgroup accuracy specifically, or even DRO generally, for Dataset Distillation. We consider a few classes of related works. In the context of DD, adversarial robustness is another popular notion of robustness, which is simply minimization with respect to the worst possible sample in the support of some perturbation on the data sample, rather than the distribution. Adversarial Robustness is more conservative than DRO, however, in some security-sensitive circumstances, this is the more appropriate framework (Wu et al., 2022). In the field of knowledge distillation, wherein a more parsimonious model, rather than dataset, is of interest, group DRO has been considered in (Vilouras et al., 2023) and (Wang et al., 2023b).
Group DRO is considered and demonstrated effectiveness for in (Sagawa et al., 2019), complemented with a regularization strategy that appears to assist in performance for small population subgroups. The coverage of disparate data distributions is also a concern in Non-IID federated learning. The work (Jiao et al., 2023) presents convergence theory for DRO in a federated setting.
@@ 3 Algorithm
Let $\mathcal{S}=\{x_{D},y_{D}\}$ be the synthetic distilled dataset, where $([x_{D}]_{i},[y_{D}]_{i})$, with $i=1,...,n_{D}$ are an individual input and label. Let $\mathcal{T}=\{x_{T},y_{T}\}$ be the training set where $([x_{T}]_{i},[y_{T}]_{j})$, $j=1,...,n_{T}$ are an individual input and label with $n_{T}\gg n_{D}$. Let $F(\theta;x,y)$ be the full loss function with parameter vector $\theta$ and data $x$ and $y$.
@@
The synthetic dataset $\mathcal{S}$ will be optimized based on an objective function that includes a risk function of the loss, averaged across a set of partitions of $\mathcal{S}$. We denote the loss function $L(\mathcal{S})$. Consider Algorithm 1. This procedure encapsulates one function evaluation of a loss with respect to the distilled data $\mathcal{S}$. Each iteration begins with subsampling the training dataset $\bar{\mathcal{T}}\subset\mathcal{T}$ and then clustering them based on their nearest synthetic data point. That is for all $i\in[|\mathcal{S}|]$ the set $\mathbf{c}_{i}$ contains the elements of $\bar{\mathcal{T}}$ that are closer to $\mathcal{S}_{i}$ than to any other point in $\mathcal{S}$.
The following criterion optimization problem is solved for $\theta$, obtaining the set of parameters that solves for minimax loss across the subgroups defined by the clusters $\{\mathbf{c}_{i}\}$
@@
$\begin{array}[]{l}\min\limits_{\theta}\left[\frac{1}{|\mathcal{S}|}\sum\limits_{i}\mathop{CVaR}{[F(\theta;\mathbf{c}_{i})]}+\max\limits_{i}\mathop{CVaR}{[F(\theta;\mathbf{c}_{i})]}\right].\end{array}$
@@
Note that rather than the expectation, we use a risk measure. We set $\alpha$ to be some tail probability. Risk measures enable one to minimize one sided tail behavior. The operator denoting the Conditional Value at Risk, $\mathop{CVaR}$, is defined, with respect to the empirical distribution of data points within $\mathbf{c}_{i}$:
@@
$\begin{array}[]{l}\mathop{CVaR}[F(\theta;c_{i})]:=-\frac{1}{\alpha}\left\{\frac{1}{|c_{i}|}\sum_{z_{t}\in c_{i}}F(\theta;z_{t})1[F(\theta;z_{t})\leq f_{\alpha}]\right.\\ \qquad\qquad\left.+f_{\alpha}(\alpha-\frac{1}{|c_{i}|}\sum_{z_{t}\in c_{i}}1[F(\theta;z_{t})\leq f_{\alpha}]\right\}\end{array}$
@@
with the quantity $f_{\alpha}$ defined as,
$f_{\alpha}:=\min\{f\in\mathbb{R}:\frac{1}{|c_{i}|}\sum_{z_{t}\in c_{i}}1[F(\theta;z_{t})\leq f_{\alpha}]\geq\alpha$
Algorithm 1 Evaluation of the Loss on the Distilled Dataset
0:   Input: Training set $\mathcal{T}$, synthetic data-set $\mathcal{S}$.
0:   0:  Execute: 0:       Subsample the training set $\bar{\mathcal{T}}\subset\mathcal{T}$ with $|\bar{\mathcal{T}}|=n_{D^{\prime}}$. 0:       Cluster $\bar{\mathcal{T}}$ by the distillation set: defining for all $t\in[n_{D^{\prime}}]$, $\mathcal{C}(z_{t})=\arg\min_{i}\|z_{t}-[\mathcal{S}]_{i}\|^{2}$ and correspondingly $\mathbf{c}_{i}=\{z_{t}\in\bar{\mathcal{T}}:\mathcal{C}(z_{t})=i\}$. 0:       Solve the following optimization problem: $\begin{array}[]{l}\min\limits_{\theta}\left[\frac{1}{|\mathcal{S}|}\sum\limits_{i}\mathop{CVaR}{[F(\theta;\mathbf{c}_{i})]}+\max\limits_{i}\mathop{CVaR}{[F(\theta;\mathbf{c}_{i})]}\right]\end{array}$ (2) to obtain $\theta^{*}$. 0:       The output $L(\mathcal{S})$ is the objective value of the optimization solution to (2).
This procedure constitutes one (noisy, due to the training subsampling) function evaluation for the dataset $\mathcal{S}$. Consider that we want to optimize $\mathcal{S}$ so as to minimize this criterion $L(\mathcal{S})$ as a function of $\mathcal{S}$. Assume that you start with some initial $\mathcal{S}$ and proceed as in Algorithm 2. To perform the full optimization, we apply a standard gradient descent approach, using some zero order approximation of the gradient of $L$ concerning the dataset $\mathcal{S}$, that is $g(\mathcal{S})\approx\nabla L(\mathcal{S})$, computed only using evaluations of $L(\mathcal{S})$.
Algorithm 2 Training the Distilled Dataset
0:   Input: Initial $\mathcal{S}$
0:   0:   Execute: 0:        For $E=1,...$ 0:          Compute: $g(\mathcal{S})\approx\nabla L(\mathcal{S})=\frac{1}{M}\sum\limits_{l=1}^{M}\frac{L(\mathcal{S}+\sigma v_{l})-L(\mathcal{S})}{\sigma}v_{l}$ with $v_{l}\sim\mathcal{N}(0,I)$ 0:          Set stepsize $s$ (diminishing stepsize rule, or stepwise decay). For instance: $s=0.1/\sqrt{1+E}$ where $E$ is the epoch 0:        Set $\mathcal{S}-sg\to\mathcal{S}$
We can use any method from (Berahas et al., 2022) to obtain a gradient estimate of $g\approx\nabla_{\mathcal{S}}L(\mathcal{S})$, here the featured procedure is Gaussian smoothing.
3.1 Analysis

### v3

@@ Abstract
Dataset distillation (DD) has emerged as a widely adopted technique for crafting a synthetic dataset that captures the essential information of a training dataset, facilitating the training of accurate neural models. Its applications span various domains, including transfer learning, federated learning, and neural architecture search. The most popular methods for constructing the synthetic data rely on matching the convergence properties of training the model with the synthetic dataset and the training dataset. However, using the empirical loss as the criterion must be thought of as auxiliary in the same sense that the training set is an approximate substitute for the population distribution, and the latter is the data of interest. Yet despite its popularity, an aspect that remains unexplored is the relationship of DD to its generalization, particularly across uncommon subgroups. That is, how can we ensure that a model trained on the synthetic dataset performs well when faced with samples from regions with low population density? Here, the representativeness and coverage of the dataset become salient over the guaranteed training error at inference. Drawing inspiration from distributionally robust optimization, we introduce an algorithm that combines clustering with the minimization of a risk measure on the loss to conduct DD. We provide a theoretical rationale for our approach and demonstrate its effective generalization and robustness across subgroups through numerical experiments.
@@ 1 Introduction
Dataset distillation (DD) is a burgeoning area of interest, involving the creation of a synthetic dataset significantly smaller than the real training set yet demonstrating comparable performance on a model (Wang et al., 2018; Sachdeva & McAuley, 2023). This practice has gained prominence in various computation-sensitive applications, providing a valuable means to efficiently construct accurate models (Gu et al., 2023b; Medvedev & D’yakonov, 2021; Xiong et al., 2023). The standard optimization objectives that are used to steer the construction of the synthetic data typically aim to foster either distributional similarity to the training set (Zhao & Bilen, 2023a; Zhao et al., 2023) or similar stochastic gradient descent (SGD) training dynamics as the original dataset (Zhao et al., 2021; Cazenavette et al., 2022). Notably, recent literature suggests that the latter category has proven more successful (Kim et al., 2022; Cazenavette et al., 2023). Intuitively, this success can be attributed to the rationale that, with considerably fewer samples, prioritizing the most relevant information for training and model building becomes more judicious.
In this paper, we aim to address two important practical concerns in DD training. First, it is essential to note that the synthetic dataset might be applied across a wide range of potential circumstances with distinctions from the training phase. Consequently, models trained on the synthetic dataset must exhibit low out-of-distribution error and strong generalization performance. This means the synthetic dataset should be designed to have certain higher-order probabilistic properties, particularly in relation to the model and loss functions. In addition, shifts in circumstances over time (some latent exogenous variables, formally) mean that the population distribution on which the model is trained may resemble a clustered sample, i.e., a subset of the population. Therefore, it is imperative that the synthetic dataset ensures good coverage across the support of the sample space.
@@
To this end, we propose to enhance the generalization performance and group coverage properties of the distilled dataset using concepts and methods in the field of Distributionally Robust Optimization (DRO) (Lin et al., 2022; Zeng & Lam, 2022; Vilouras et al., 2023). Within DRO, one solves a bilevel optimization problem where the objective is to minimize the loss over a data distribution, subject to the constraint that this distribution is a specific probabilistic distance away from the population distribution:
$\min\limits_{\theta\in\mathbb{R}^{d}}\max\limits_{\mathcal{Q}\in\mathbb{Q}},\,F(\theta,\mathcal{Q}),\,\mathbb{Q}=\{\mathcal{Q}:\mathcal{D}(\mathcal{Q}||\mathcal{P})\leq\Delta\}$ (1)
@@
where $F$ is the loss function, $\theta$ is the parameter, $\mathbb{Q}$ is the distributional uncertainty set, $\mathcal{D}$ is an f-divergence and $\mathcal{P}$ is the population distribution.
@@
We now present the two desiderata of DD to which we aim to apply the concepts of DRO. First, our goal is to construct a synthetic dataset that is a quality representative of the underlying population distribution. Thus, the consideration of generalization is paramount. Xu & Mannor (2012) present arguments indicating the theoretical equivalence of testing to training error as robustness to perturbations of the data distribution, and the generalization accuracy. Second, a conventional maximum likelihood classifier inherently assigns higher weights to sample ranges with larger prior distributions than to those with lower overall probability density. This inherent bias poses a fundamental risk of underperformance in small population subgroups. Furthermore, in online inference, it is common to observe alterations in the overall prior distribution, particularly in terms of subgroups. DRO has been conjectured and observed to mitigate this issue, both generically (Duchi & Namkoong, 2021) and with explicit quantification of subgroups in the distributional uncertainty set (Oren et al., 2019), where subgroups of topics are considered for training language models.
Figure 1: A visual representation of the robust inference task involves the partial partitioning of the population distribution, that is $\mathcal{X}$ across subgroups $\{\mathbf{c}_{i}\}$. A classifier is considered robust when it demonstrates high performance across the subgroups. As a practical hypothetical example of online learning, at a particular time, a steady stream of samples from $\mathcal{X}|\mathbf{c}_{3}$ may appear to the classifier. Note that in this case the region of sample space defined by this subgroup is geometrically small, and we can consider that it has a low overall prior density. If this subgroup’s behavior is particularly anomalous, a model and any associated distilled dataset trained only on minimizing empirical risk may perform poorly on this subgroup.
Moreover, this permits flexibility as far as easily modularized methods to specific concerns. Indeed, domain expertise in regards to the properties of the population distribution has been shown to assist in generalization– (Hu et al., 2018) indicates that a more precisely defined distributional uncertainty set, focusing on specific collections of probability distributions from subsamples of the population rather than encompassing all possibilities in a ball, yields improved generalization performance.
@@
Inspired by these concepts, this paper introduces a Double-DRO-based DD procedure designed specifically to address these two concerns. Fundamentally we employ DRO at two levels to promote different advantages that we shall witness in the experimental results. 1) We leverage the loss across clusters of the latent variables to ensure group robustness. 2) We approximately solve a DRO problem internally in order to promote better within-group generalization, using a risk measure as a proxy for a DRO solution.
@@
Our paper continues below as follows. Next, we present our algorithm, we analyze the optimization problem defining the notion of generalization of interest, and argue how our method effectively targets this criterion in Section 2. We give the mathematical analysis in Section 3. In Section 4 we present numerical results validating the superior overall and subgroup generalization performance on standard machine learning benchmarks. Finally, we conclude this work in Section 5. Related works are reviewed in the Appendix in Section E.
@@ 2 Algorithm
In this Section we describe our main procedure, in which we incorporate two techniques informed by DRO in order to improve both generic generalization performance and group distributional coverage.
@@
Let $\mathcal{S}=\{x_{S},y_{S}\}$ be the synthetic distilled dataset, where $([x_{S}]_{i},[y_{S}]_{i})$ with $i=1,...,|\mathcal{S}|$ is an individual input and label pair. Let $\mathcal{T}=\{x_{T},y_{T}\}$ be the real training set where $([x_{T}]_{i},[y_{T}]_{j})$, $j=1,...,|\mathcal{T}|$ is an individual input and label pair with $|\mathcal{T}|\gg|\mathcal{S}|$. The target of dataset distillation is to optimize the synthetic dataset $\mathcal{S}$ so that a network with parameter $\theta$ can achieve similar performance compared with that attained on $\mathcal{T}$. For simplicity, we define the input and label pair as $z=(x,y)$. And $F(\theta;z)$ is the full loss function to optimize parameter $\theta$ on data $z$, which is typically cross entropy loss in classification tasks.
Matching-based dataset distillation methods usually involve imitating certain training characteristics of the real training set, such as training gradients, feature distribution, training trajectory, etc (Zhao et al., 2021; Zhao & Bilen, 2023a; Cazenavette et al., 2022). Take gradient matching as an example. At each iteration, the training gradient is extracted for the synthetic data and the real data, based on the same network. The gradient difference is set as the objective of optimizing synthetic data. Considering that diverse supervision gradients from different training stages should be provided to enhance the consistency throughout a model training process, the network is simultaneously updated with the synthetic data optimization. Previous methods utilize the same loss function $F(\theta;z)$ to update the network. However, in this work, we propose to incorporate DRO at this stage for more robust network optimization as well as distillation supervision.
Specifically, we wish to solve a DRO locally in order to promote solutions that are particularly suitable for generalization. However, as the dimensionality and desired training size in our settings of interest present intractability for DRO, instead of solving the full DRO we use a proxy in the form of a Conditional Value at Risk (CVaR), which we define below. In the theoretical analysis, we describe the known correspondence between CVaR and DRO.
@@
In our algorithm, each iteration begins with subsampling the training dataset $\bar{\mathcal{T}}\subset\mathcal{T}$ and then clustering them based on their nearest synthetic data point. That is for all $i\in[|\mathcal{S}|]$ the set $\mathbf{c}_{i}$ contains the elements of $\bar{\mathcal{T}}$ that are closer to $z_{i}$ than to any other point in $\mathcal{S}$. The following criterion optimization problem is solved for $\theta$, obtaining the set of parameters that solves for minimax loss across the subgroups defined by the clusters $\{\mathbf{c}_{i}\}$
@@
$\begin{array}[]{l}\min\limits_{\theta}\left[\frac{1}{|\mathcal{S}|}\sum\limits_{i}\texttt{CVaR}{[F(\theta;\mathbf{c}_{i})]}+\max\limits_{i}\texttt{CVaR}{[F(\theta;\mathbf{c}_{i})]}\right].\end{array}$ (2)
By considering the clusters $\mathbf{c}_{i}$ separately, especially including the $\max\limits_{i}$ term in the objective, we robustify the performance with respect to the different clusters. The synthetic dataset is meant to represent the training set, and with clustering, we ensure the entire support of the training distribution is represented by some sample $z_{i}\in\mathcal{S}$. This yields our effort to ensure the group distributional robustness. We intend that both the worst case and the average performance across the sub-groups are minimized.
@@
In addition, to promote generalization broadly speaking, note that rather than the expectation, we use a risk measure, as far as statistically agglomerating the loss over the training data. We set $\alpha$ to be some tail probability. Risk measures enable one to minimize one-sided tail behavior. The operator denoting the Conditional Value at Risk, CVaR, is defined, with respect to the empirical distribution of data points within $\mathbf{c}_{i}$:
@@
$\begin{array}[]{l}\texttt{CVaR}[F(\theta;\mathbf{c}_{i})]:=-\frac{1}{\alpha}\left\{\frac{1}{|\mathbf{c}_{i}|}\sum_{z_{t}\in\mathbf{c}_{i}}F(\theta;z_{t})\mathbf{1}[F(\theta;z_{t})\leq f_{\alpha}]\right.\\ \qquad\qquad\left.+f_{\alpha}(\alpha-\frac{1}{|\mathbf{c}_{i}|}\sum_{z_{t}\in\mathbf{c}_{i}}1[F(\theta;z_{t})\leq f_{\alpha}])\right\}\end{array}$ (3)
@@
with the quantity $f_{\alpha}$ defined as, with $|\mathbf{c}_{i}|$ denoting the size of the training set in cluster $\mathbf{c}_{i}$,
$f_{\alpha}:=\min\left\{f\in\mathbb{R}:\frac{1}{|\mathbf{c}_{i}|}\sum_{z_{t}\in\mathbf{c}_{i}}\mathbf{1}[F(\theta;z_{t})\leq f_{\alpha}]\geq\alpha\right\}$ (4)
Algorithm 1 Robust Dataset Distillation
0:   Input: Real training set $\mathcal{T}$, synthetic set $\mathcal{S}$, network with parameter $\theta$, distilling objective $L(\mathcal{S})$.
0:   0:  Execute: 0:      While not converged: 0:       Subsample the training set $\bar{\mathcal{T}}\subset\mathcal{T}$. 0:       Cluster $\bar{\mathcal{T}}$ by the distillation set, i.e. define, for all $t\in[|\bar{\mathcal{T}}|]$: 0:         $\mathcal{C}(z_{t})=\arg\min_{i}\|z_{t}-[\mathcal{S}]_{i}\|^{2}$ and 0:         $\mathbf{c}_{i}=\{z_{t}\in\bar{\mathcal{T}}:\mathcal{C}(z_{t})=i\}$. 0:       Solve the optimization problem in equation 2 to obtain $\theta^{*}$. 0:       Optimize synthetic set $\mathcal{S}$ with $L(\mathcal{S})$ based on the optimized parameter $\theta^{*}$.
The choice of a CVaR weighing of the loss informed by DRO-associated theoretical work aims to improve test error accuracy rather than merely minimizing training error. Additionally, by clustering and weighting the balanced loss, Group DRO is applied across the connected components of the sample space distribution, which defines the population.
After updating the network with Group DRO, the synthetic dataset $\mathcal{S}$ will be optimized based on matching metrics, denoted by $L(\mathcal{S})$, where existing matching-based DD methods can be broadly plugged in. This makes the algorithm modular to any choice of DD procedure. The algorithm is shown in Alg. 1. As $L(\mathcal{S})$ involves a risk function of the loss $F(\theta;z)$, averaged across a set of partitions of $\mathcal{S}$, the supervision helps enhance the distributional robustness of the distilling process. Thereby, the distilled dataset can obtain superior generalization performance across different domain shifts, as well as better group coverage for more practical usefulness. We proceed to iterate between an update to the parameter based on solving equation 2 and an update to the synthetic dataset $\mathcal{S}$ with our tailored DD procedure until the procedure reaches a fixed point, wherein the two do not significantly change.
3 Analysis
In this Section we describe the formal optimization problem being solved as well as its convergence and statistical properties. The focus will be on justifying why we expect the procedures defined in the Algorithm to improve (group) generalization, through foundational results in DRO theory.